### v2

# Did I Vet You Before? Assessing the Chrome Web Store Vetting Process through Browser Extension Similarity

Abstract
Web browsers, particularly Google Chrome and other Chromium-based browsers, have grown in popularity over the past decade, with browser extensions becoming an integral part of their ecosystem. These extensions can customize and enhance the user experience, providing functionality that ranges from ad blockers to, more recently, AI assistants. Given the ever-increasing importance of web browsers, distribution marketplaces for extensions play a key role in keeping users safe by vetting submissions that display abusive or malicious behavior. In this paper, we characterize the prevalence of malware and other infringing extensions in the Chrome Web Store (CWS), the largest distribution platform for this type of software. To do so, we introduce SimExt, a novel methodology for detecting similarly behaving extensions that leverages static and dynamic analysis, Natural Language Processing (NLP) and vector embeddings. Our study reveals significant gaps in the CWS vetting process, as 86% of infringing extensions are extremely similar to previously vetted items, and these extensions take months or even years to be removed. By characterizing the top kinds of infringing extension, we find that 83% are New Tab Extensions (NTEs) and raise some concerns about the consistency of the vetting labels assigned by CWS analysts. Our study also reveals that only 1% of malware extensions flagged by the CWS are detected as malicious by anti-malware engines, indicating a concerning gap between the threat landscape seen by CWS moderators and the detection capabilities of the threat intelligence community.
I Introduction
Over the past decade, web browsers have become an indispensable tool for both desktop and mobile users. As traditional applications migrate to web-based services, users spend more time in their browsers, increasing the need for extended and more customization features. Chromium-based browsers—notably, Google Chrome—allow this customization through browser extensions: third-party programs built with web technologies that interact with the browser to add or modify features, and enhance the user experience. Some popular types of extensions include ad blockers, grammar checkers, password managers, and, more recently, AI assistants.
The undisputed marketplace for browser extensions for Chromium browsers is the Chrome Web Store (CWS), Google’s official marketplace which launched back in 2011 [30]. The CWS plays a vital role in the browser ecosystem as many users inherently trust the extensions included on its catalog. Similar to the case of the Google Play Store for Android apps, the CWS implements vetting processes to detect and mitigate the risk of harmful and deceptive extensions being distributed through the official store. Extension publishers must adhere to the CWS Program Policies, which govern what content can be uploaded and distributed [23]. These policies forbid not only the publication of harmful content, but also mandate that extensions must disclose their behavior, implement a minimum of functionality, and not duplicate existing content, among other requirements. If an infringing extension (i.e., one that does not abide by the CWS policies) is published, Google may delist it at its discretion during the vetting process and, in some cases, even remove it from browsers that have it installed.
In the past, the CWS has been under scrutiny from security experts and researchers. Various reports and research studies identified infringing content being distributed on the CWS, focusing mostly on potentially harmful extensions [46, 61, 53] and in methods to detect their presence [29, 28, 1, 48, 51]. While there is evidence of Google proactively vetting extensions from its official store [28], the effectiveness of the vetting process at detecting infringing content remains unexplored. Recent research has focused on characterizing the content of the store and its trends [26], including the presence of extensions with vulnerabilities or a similar code base. However, a knowledge gap remains when it comes to understanding Google’s policing of the platform. Characterizing the challenges that the CWS vetting process faces in practice can help identify its pitfalls and improve its effectiveness. To fill this gap, this paper aims to answer the following research questions:
  • RQ1.
    How effective is the CWS vetting process at detecting and rapidly removing infringing extensions?
  • RQ2.
    What are the main kinds and the key features of infringing extensions uploaded to the CWS?
  • RQ3.
    What malware families target the CWS and what threat intelligence is available about them?
To answer these questions, we develop the scalable data collection and hybrid analysis pipeline for browser extensions shown in Figure 1. We compile a dataset of 366,617 extensions downloaded from the CWS during 4 years. We enrich this dataset with vetting labels assigned by Google moderators to removed CWS items. These labels indicate which extensions have been taken down from the store and for what reason, thus providing ground truth about the vetting process. To the best of our knowledge, this is the first work that studies the CWS vetting process using ground truth about removed content. We then design and develop SimExt, a novel methodology for finding similarly behaving extensions that leverages $(i)$ static and dynamic analysis for extracting syntactic and behavioral features; and $(ii)$ Natural Language Processing (NLP) and vector embeddings for computing similarity. To address the existing technical gap in feature extraction from browser extensions, we develop and release Fakeium [42], an open source dynamic analysis sandbox that follows an implementation inspired by concolic execution to elicit behaviors and overcome the limitations of static analysis and fuzzy-hash-based methods [26]. Fakeium is very scalable and can process an extension in mere seconds, which makes it suitable for large-scale measurements such as the one conducted in this work. To generate the vector embeddings, we use a novel approach based on Zero-Short Learning (ZSL) which, unlike previous work, does not require retraining the model when new malicious extensions are uploaded to the store [28, 1, 48, 51].
Using SimExt, we perform a large-scale behavioral cluster analysis to find published extensions that are very similar to vetted ones, as these are natural candidates for being infringing items that were overlooked by the vetting process. Equipped with this pipeline, we make the following novel contributions about the CWS vetting process and its effectiveness:
  1. 1.
    We find 17,021 potentially infringing extensions that are still published on the CWS. Our pipeline flags these extensions as extremely similar to known infringing extensions previously taken down from the store. Interestingly, 86% of these infringing items are republished extensions, i.e., extensions that are either identical or extremely similar to items that were previously taken down, suggesting that the CWS vetting process does not adequately learn from experience. We also find that 11% of the infringing extensions that remain published at the end of our crawl come from repeat offenders (publishers with known vetted items), suggesting that the CWS does not suspend these accounts according to their own policies. We also conduct a survival analysis for infringing extensions. Our results reveal that the process of identifying infringing extensions is excessively slow, taking months or even years from the publication of an infringing version to its removal from the store.
  2. 2.
    To validate our findings, we analyze extensions from the top infringing clusters with the most extensions and manually assigned them a type or kind of content. We find that 83% of the extensions in these clusters are New Tab Extensions (NTEs), which are items that override the webpage that loads when a new tab is opened. The cluster analysis also reveals the presence of spam and no content extensions, both vetted and still published in the CWS. In order to better understand the labeling process, we explore the relationship between the CWS vetting labels and the behavior of an extension. Our analysis finds that labels are somewhat inconsistent and unrelated to features or behaviors, especially in the case of NTEs.
  3. 3.
    We scan the 5,647 extensions labeled as malware by the CWS in VirusTotal and compare the results. Remarkably, 95% of malware-labeled samples were not indexed by VirusTotal and 3% were known but have zero detections. This finding suggests that the majority of malicious extensions detected in the CWS are unknown to the threat intelligence community, and that for others there is a significant discrepancy between the criteria of CWS moderators and the detection capabilities of malware detection engines. Analysis of the malicious extensions for which we obtain detection reports reveals that the malware lineage analysis conducted by security vendors is very poor, with the family label missing in most cases and being very general in others.
Disclosure. We reported to Google a subset of 180 infringing extensions found by our pipeline. All of them were manually reviewed to confirm the presence of a policy violation. Most importantly, this set includes 23 no content extensions published from corporate Google accounts, all of which were removed from the store shortly after our report. As of this writing, we have not received an official response.
Tool and Research Artifacts. We release the source code of Fakeium, an instrumented JavaScript sandbox that can be used to extract API calls from browser extensions [42]. We also provide the list of extension pairs used as ground truth for evaluation, and metadata and high-dimensional embeddings for infringing extensions found by our pipeline at https://zenodo.org/records/10977708.
II Background
Figure 1: Data collection and analysis methodology pipeline.
This section provides background on browser extensions and the CWS publishing ecosystem.
II-A Browser Extensions
Browser extensions are third-party programs that enhance or modify the functionality of the browser, with ad blockers being one of their most recognizable and widely-used examples. Extensions are written using a combination of HTML, CSS and JavaScript, and may contain other resources commonly found in web applications, such as images or fonts [39]. Extensions are packaged as CRX files, bundling all the source code and assets needed to run them into a single file, thus easing distribution. CRX packages are regular ZIP archives with an additional header that ensures the integrity and authenticity of its contents [49].
As in the case of other popular platforms like Android, all extensions have a mandatory manifest.json file that dictates their functionality and behavior. This manifest is a JSON document that specifies basic metadata about the extension, like its name and version. Optionally, it may also declare permissions, service workers (which run code in the background), content scripts (which run code in open tabs), and overridden pages (e.g., to modify the default URL when opening a new tab), among other elements [10]. Two popular types of extensions are themes and New Tab Extensions (NTEs). A theme is a type of extension that changes the way the browser looks and does not contain any source code, i.e., all theme configuration is in the manifest itself. An NTE is an extension that changes the default webpage that will load every time a new tab is opened.
Extensions communicate with the browser using Web APIs [40] and Extension APIs [21]. These APIs provide extensions with a broad range of powerful capabilities, allowing them to access and manipulate the Document Object Model (DOM) of webpages, and to access geolocation data, browsing history, and other sensitive information stored in the browser. Some of these APIs are restricted by default for security reasons, and can only be accessed by declaring the appropriate permissions in the manifest.
II-B CWS Program Policies
CRX packages can be published on the Chrome Web Store (CWS), Google’s official distribution platform for browser extensions that run on Google Chrome [22]. Users of other Chromium-based browsers, such as Microsoft Edge, Brave, or Opera, can also install extensions from this source as they run the same compatible engine. In fact, Brave even recommends that users download extensions from this platform in the absence of their own official store [2]. As such, policing the content that is uploaded to the CWS is of great importance because of the impact that harmful extensions can have given the large audience of the store.
The CWS requires all developers to comply with a set of policies that include security and privacy guidelines. These policies prohibit the publication of extensions with a malicious purpose and extensions that facilitate unauthorized access to restricted or copyrighted content [17]. They also ask developers not to obfuscate code or conceal functionality, to request the minimum set of permissions needed, and forbid submitting multiple extensions that provide duplicate or very little functionality [16, 15, 18]. According to Google’s policies, repeated violations of these policies will result in the suspension of the publisher account [19].
III Data Collection
We implemented a purpose-specific crawler to download the resources shown in Figure 1, namely CRX packages and their associated metadata, from Google’s Chrome Web Store [22]. Our crawling strategy consists of three steps. First, we prepare an initial set of extensions from the store’s sitemap, the reels of popular items from each category, and the featured extensions from the home page. We then visit each extension page, extract metadata (e.g., name, description, number of installs, last modification date) and download its associated assets (i.e., CRX package, icon and tile). Finally, we look for non-visited extensions appearing as “related” or “recommended” in the store page of each visited extension and add them to the queue of items to crawl. Previous efforts demonstrate the effectiveness of this approach [9, 31], which takes inspiration from strategies successfully used to crawl the Google Play Store [65, 50].
We run the crawler daily to detect new and removed extensions, and keep track of changes being pushed to the store such as publishers uploading an updated version of an existing extension. For simplicity, we set the visitor’s country code to the United States when crawling the store. Although developers can opt-out of publishing extensions in certain countries or even entire regions [14], this decision should not limit our coverage as we can still download any non-private CRX package regardless, and because the used sitemap contains publicly listed extensions worldwide. We only run our data collection pipeline once a day and rely on the HTTP caching headers provided by CWS to responsibly limit our impact on Google’s servers. We note that in October 2023, we transitioned to crawling the new version of the CWS, coinciding with its first public release [27]. We made this decision in preparation for the announced retirement of the previous version in January 2024. While we experienced no issues during this migration, we also kept crawling the legacy store for a few more weeks in the event something went wrong.
In addition to crawling CRX packages and store metadata, we also fetch vetting labels issued by Google from a separate endpoint. These labels indicate whether an extension was taken down from the CWS instead of willingly being unpublished by its developer, as well as the reason behind the takedown (e.g., flagged as malware), if any. Vetting labels are used by Google Chrome since version 117 to protect users of vetted extensions by proactively disabling malware and displaying warnings in the browser extensions page [7]. As far as we can tell, this is the first study to crawl and analyze this data.
III-A Dataset
We crawled the CWS for 4 years from March 2020 to March 2024. We compiled a historical dataset of 366,617 extensions, containing 902,532 different updates (or extension versions) that were available for download at some point in the store. Some of the extensions in our datasets were published as early as December 2009.111 While the CWS was officially launched in 2011, it existed prior to that date as evidenced by https://narkive.com/bEuDVJqN Of all these versions, we lack the CRX package for 28,405 paid extensions that did not change their pricing model after payments were deprecated in the CWS in mid-2020 [13]. Our dataset also contains 53,132 browser themes, which are distributed in the same form as extensions but do not have the ability to execute code. To the best of our knowledge, this is the most comprehensive dataset of browser extensions to-date, with over 156k more extensions than the previous large-scale work that crawled the CWS [48].
Figure 2: Daily volume of published extensions (solid, left axis) and accumulated dataset size (dashed, right axis).
Figure 2 shows the change over time of the number of currently published extensions in the CWS (solid line) versus the total size of our dataset (dashed line). We observe that our dataset grows by an average of 106 new extensions per day, with this rate remaining fairly consistent throughout the 4 years of our study. Yet, the number of published extensions in the store has decreased by 17% since March 2020, experiencing a noticeable drop in December 2020 and following a downward trend since then. This contrasts with the 74% increase of our dataset in the same time period. Without further analysis of the dataset, this growth suggests a very dynamic ecosystem with a high rotation of extensions entering and leaving the store. This rotation could either be attributed to developers deliberately publishing and shortly thereafter unpublishing their extensions, or Google taking down a large number of malware and other abusive content as part of its vetting process.
Vetting Labels. Using the vetting labels that we crawl for removed extensions, we can infer whether they were unpublished by the developer or taken down by Google. Vetted extensions have a type of either Malware, Policy Violation, or Minor Policy Violation, whereas items that were unpublished at the request of their developers have a type of None.
Figure 3 shows the ratio of extension labels found in our dataset at a given point in time since we first started crawling. As of March 2024, 52% of all the extensions we collected have been removed from the store. More interestingly, 79% of the removed items are infringing extensions that were taken down. This preliminary characterization of the dataset motivates us to delve deeper into the dynamics of extension removals and the vetting process of the CWS.
Figure 3: Daily ratio of vetting labels provided by CWS.
IV SimExt: A Methodology for Finding Similar Extensions
While developing a fully automated vetting process is very challenging, having an automatic approach to detect similar extensions at scale is valuable to assist analysts, as it can be used to find items that resemble previously vetted ones or clones to currently published content. Our key working hypothesis is that if an extension is taken down from the store because it violates a program policy, then other extensions that are highly similar are candidates to be vetted as well. Even in those cases where the behavior of an extension is not necessary harmful, finding clones is still valuable given the CWS ban on repetitive content [15]. We acknowledge that legitimate extensions cloned by malicious actors are an exception to this rule. For example, the vetting of malicious extensions reusing MetaMask’s codebase should not result in MetaMask being taken down. We consider that these cases should be manually filtered out, and that a tool that provides detailed similarity reports between two extensions can be extremely helpful to assist the store vetting process. Section VI provides examples of published infringing extensions found by our tool to support this claim.
Since there is currently no tool that identifies similar extensions for a given input item, we designed and built SimExt, our own state-of-the-art solution for querying browser extensions based on similarity. We design it as a multi-step pipeline to automatically extract meaningful behavioral features from all extensions in our dataset and then group them into clusters based on their similarity. This approach is different from previous work that classifies extensions as either benign or malicious [28, 1, 48]. Moreover, to the best of our knowledge, our work is the first one to leverage Natural Language Processing (NLP) and vector embeddings to effectively and accurately find similar extensions at scale. Figure 1 provides a summary of the pipeline steps described below.
IV-A Feature Extraction
SimExt uses both static and dynamic analysis to extract manifest properties and API calls to determine the structure and behavior of an extension.
Manifest. We flatten the properties found in the extension manifest to convert them from a nested JSON object to key-value pairs. We exclude the author, name, short_name, description, version, key and update_url properties as they typically do not provide reliable information about the behavior of an extension and, in the worst case, could help the attacker to trick the model into giving a different embedding to the extension. We also identify and enumerate all code entrypoints found in the manifest. Per Google’s documentation [10], we consider any of the following items to be valid entrypoints: $(i)$ content scripts; $(ii)$ background pages or service workers; $(iii)$ popup pages, i.e., browser actions; and $(iv)$ overridden, DevTools, side panel, or options pages.
Static Analysis. We statically analyze the JavaScript code of entrypoints. We use esbuild [63] for module resolution, bundling, tree shaking (i.e., dead code removal) and minification. We then use Babel [43], a well-maintained and widely used tool in the modern web ecosystem, to generate the Abstract Syntax Tree (AST) of the resulting bundle. Although previous work has relied on Esprima [25] for this purpose [9, 48, 58, 59, 8], we decided to use a more modern alternative given that Esprima has not received any major updates since 2018. Lastly, we traverse the AST to extract the list of API calls as relevant features to model the behavior of the extension. We define an API call as an invocation to an Extension API [21] from the chrome or browser object, or to a Web API [40] within the navigator object. In particular, we use the Babel Traverse module and the ReferencedIdentifier and BinaryExpression nodes to find the names of global variables (e.g., with no bindings except for the global scope) being invoked in a program.
Dynamic Analysis. To overcome the limitations of static analysis and detect API calls we might otherwise miss (e.g., code being executed in a call to eval), we complement our feature extraction pipeline with a dynamic analysis stage. We develop Fakeium, our own sandbox environment based on isolated-vm [32] to safely run untrusted JavaScript code. Fakeium intentionally lacks the webpage DOM and Web and Browser APIs. Instead, it works by mocking any objects accessed by the extension at runtime to prevent it from crashing for as long as possible. This mocking is performed by hijacking the global scope using a Proxy object [38]. We monitor calls to mocked objects and log API calls as in static analysis. To ensure that we cover as many code paths as possible, we follow an approach inspired by concolic execution, where we invoke all callbacks and functions found during the analysis, whilst still respecting the original conditional predicates. This approach enables Fakeium to run without the need for an automation tool or monkey to trigger behaviors. While not as accurate as running the extension in an actual web browser, this approach is more scalable, taking just a few seconds instead of minutes to analyze an extension. In addition, this environment relies on the same JavaScript engine that Chromium uses [60], guaranteeing accurate output for code snippets that depend only on the ECMAScript specification (i.e., without invoking APIs that will be mocked by our sandbox). Fakeium is released under the MIT license and distributed as a Node.js package [42].
IV-B Vectorization
We apply Natural Language Processing (NLP) to generate a vector embedding for each extension based on its features. First, we serialize manifest properties and API calls into text sentences that are sorted alphabetically and concatenated with a semicolon to produce documents. Specifically, we remove all punctuation and normalize word casing for manifest keys and API calls, and perform a laxer transformation for manifest values. We merge similar manifest keys together and limit the number of values per group to avoid extremely long documents. Appendix A provides an example of these feature serialization techniques for reference.
Inspired by Zero-Shot Learning (ZSL) [52], we use Sentence-Transformers [54] with the “all-distilroberta-v1” variant of the DistilRoBERTa model222 See https://huggingface.co/sentence-transformers/all-distilroberta-v1 to compute 768-dimensional embeddings of the generated documents. DistilRoBERTa is a general purpose model based on the BERT framework that was trained on a large and diverse corpus of English texts [34]. It uses knowledge distillation to improve performance while retaining a high accuracy [56]. We found this model to be a good fit for our needs as it achieves a good balance between sentence performance and encoding speed.333 See comparison at https://www.sbert.net/docs/pretrained_models.html Using the default tokenizer provided with the model, the average length of our documents is 300 tokens, with a standard deviation of 228 tokens. While some inputs are truncated after exceeding the maximum sequence length of 512 tokens set by the model, this number is fairly small as 86% of the documents fit within the limit.
IV-C Clustering
We employ density-based clustering to group together extensions with a very similar behavior based on their embeddings, and to filter out outliers that do not belong to any cluster. Prior to the clustering, we use Principal Component Analysis (PCA) to reduce the dimensionality of the standardized embeddings for performance while retaining 95% of the amount of variance. This results in PCA deciding to keep 161 components. Since DistilRoBERTa already outputs normalized embeddings, we skip the standardization step. Then, we use Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) [36]. We choose HDBSCAN over other clustering algorithms like K-Means as it does not require knowing the number of clusters beforehand. We pick a minimum cluster size of 5 data points and lower the minimum samples parameter to 2. This configuration should discard points with low stability (i.e., outliers) while still allowing us to create clusters in less dense areas.
Visualization. To facilitate the visualization of the extensions in our dataset and their similarities, we use Uniform Manifold Approximation and Projection (UMAP) [37] to project the high-dimensional space formed by the standardized extension embeddings onto a 2D plane. This step does not affect the results of the clustering process, as we rely solely on HDBSCAN for assigning cluster IDs to extensions.
IV-D Evaluation
To validate the performance of SimExt, we use a dataset of similar and different extensions and check whether the pipeline correctly puts similar pairs in the same cluster and different pairs in different clusters.
Ground Truth. Given the lack of previous work in this area, we could not find a dataset of similar CRX files, nor a tool to facilitate this process. Therefore, we create our own ground truth by listing popular extensions in the CWS and using the embeddings generated by our pipeline to find similar versions of these items. Additionally, we take advantage of developers publishing the same extension across stores to add them to the list of similar pairs. That is, we manually look for samples of featured CWS extensions that are also published on Edge Add-ons [41] or Opera add-ons [45], as we expect those to be identical. We build pairs of different extensions by looking for items that offer the same functionality but implement it differently. For example, uBlock Origin and Privacy Badger make a pair of different ad blockers.
We consider two items to be similar if they meet all of the following criteria: $(i)$ their manifests have an overlap of $\geq$90of their keys; $(ii)$ their manifests have unique values in common, such as file paths or localized strings; $(iii)$ their file trees have $\geq$50of paths in common, excluding localized messages.json files; and $(iv)$ they share source code files that are identical after beautification, excluding third-party libraries.
For a fairer evaluation, we complement the ground truth with manually-reviewed pairs of similar and different extensions taken from random clusters obtained after building the pipeline. We provide the resulting list of pairs as an artifact for reproducibility (see Section I).
Results. Using the process described above, we managed to put together a ground truth dataset of 110 pairs of similar extension versions, and another 110 pairs of different ones (220 unique pairs in total). When validating against the comprised ground truth, SimExt shows an accuracy of 83.2%, with a precision of 85.4% and a recall of 80.0%. We consider these results adequate for the task: as the first work to use embeddings for clustering browser extensions, the high accuracy obtained gives us confidence in the proposed novel methodology and positions us well to conduct the first examination of the CWS vetting process.
V Effectiveness of the Vetting Process
This section examines the effectiveness of the CWS vetting process at identifying content that fails to comply with store policies and the time that offending items remain published before they are eventually taken down (RQ1). To answer this question, we take the most recent version for each extension (including removed extensions) as of the end of our crawl in March 2024.
V-A Infringing Extensions Detection
As discussed in Section IV, we assume that extensions that are very similar to vetted extensions are most likely in violation of store policies and should likely be taken down. We acknowledge that this is a strong assumption, and we validate it in Section VI by performing a manual analysis of the clusters and providing compelling examples of nearly identical extensions found by SimExt. To benchmark how good the CWS is at finding these potentially infringing extensions, we quantify how many extensions that are still published as of the latest snapshot of our dataset belong to a vetted cluster. We define an infringing cluster as a set of 2 or more similarly behaving extensions in which at least one of them has been taken down by Google (i.e., has been vetted) according to the the labels they provide, as presented in Section III.
(a) Removed extensions
(b) Published extensions
Figure 4: Scatter plots of extensions belonging to an infringing cluster. Taken down (vetted) extensions in red, unpublished by the developer in green, and still published in blue.
Figure 4 shows two-dimensional projections of the embeddings of extensions found in infringing clusters. In total, our tool finds 6,444 infringing clusters comprised of 100,957 extensions. These clusters contain 72,699 extensions that were taken down (72%), and an additional smaller—yet significant—set of 11,237 unpublished extensions (11%). According to CWS labels, the aforementioned were voluntarily unpublished by their developers without Google raising a policy violation despite being similar to vetted extensions. In Section VIII we discuss why we believe publishers might prefer to intentionally remove their extensions over receiving a takedown by Google. Lastly, our pipeline finds 17,021 published extensions that account for the remaining 17% of infringing clusters. Overall, these numbers are reasonably good and show that the CWS has some capability to detect infringing extensions. Nevertheless, we reiterate that the number of published infringing items is an estimate derived from our pipeline and the actual amount may vary.
Impact. We measure the impact of infringing extensions based on the number of active user installs, using the most recent count available. Removed extensions have a cumulative sum of 802M users, whereas published extensions have 172M users in total. The mean for removed items stands at 9.6k±$\pm$108k users, indicating substantial variability, with 25%, 50%, and 75% of installs lying below 1, 12, and 146, respectively. For published extensions, this distribution is slightly higher, with users falling beneath 1, 17, and 235 for the same ranges above. As such, we find that the number of users does not meaningfully change between removed and published but potentially infringing extensions. This suggests that Google removes items regardless of their popularity.
Detection Rate. To assess how effective the CWS is at detecting infringing extensions across clusters of different size, we define the the detection rate for a cluster as the number of vetted extensions in the cluster divided by its size. Ideally, a perfect similarity tool with a flawless vetting process should provide 100% rate for all infringing clusters, implying that all policy-violating items are taken down by the store operator. Instead, our tool finds that only 1,397 clusters (22%) have a perfect rate. On average, infringing clusters have a detection rate of 0.59, widely ranging from 0.33 to 0.88 from the lower to the upper quartile.
Republished Extensions. In terms of effectiveness at finding infringing extensions, we also examine whether the CWS learns from experience and blocks the publication of extensions that are similar to previously vetted ones. To answer this, we take the earliest vetted extension from each infringing cluster and count the number of republished extensions, i.e., similar items that were published at a later date. According to our pipeline, 92% of infringing clusters contain extensions that were published after the fact. Furthermore, republished extensions make up to 86% of all infringing extensions, or 86k items in total. These results ultimately suggest that CWS lacks rules for blocking extensions similar to known vetted items.
Repeat Offenders. We find evidence of 6,156 repeat offenders, or publishers with multiple vetted extensions. Up to 11% of these publishers (663 accounts) have at least one published infringing extension as of our last date of crawling. Put another way, 11% of the published infringing extensions (1,896 items) come from repeat offenders. Another takeaway is that these accounts were not banned from the store after repeatedly uploading confirmed infringing extensions. According to the CWS, “repeated violations” of the Program Policies will result in the suspension of the developer account [19]. The existence of known repeat offenders with published extensions is a strong indication that this policy is not being properly enforced.
Takeaway. We estimate that the CWS fails to detect 17% of extensions which are similar to previously vetted items, suggesting it lacks the ability to query extensions based on behavioral similarity. As 86% of infringing items are republished extensions, using similarity search will greatly improve the vetting process. Repeat offenders are also a concern, as 11% of published infringing extensions come from accounts with known violations that have not been suspended.
V-B Survival Analysis
We perform a survival analysis [4] on infringing extensions to assess how rapidly the CWS removes policy-violating content. To determine how long these extensions remain published in the store, we measure the lifetime or duration in days between the release date of the infringing extension version and its “death” (removal date). For published infringing extensions—for which we do not observe this event—, we take the latest date of our crawling as the removal date, and label them as right-censored observations.
The lifetime distribution for vetted extensions provides some worrying results: 59% of vetted extensions stay in the CWS for more than a year, with 23 items even surpassing a decade before being taken down. Only 13% of these extensions are vetted in less than a month, and merely 4% within a week. These figures suggest that it takes several months or even years for Google to remove infringing extensions from the CWS. During this time, Chromium users can download and install these items without any notice.
Figure 5: Kaplan–Meier curves of the survival of infringing extensions by vetting label.
To study the likelihood of survival for all infringing items, we use the lifelines Python library [6] to plot Kaplan–Meier curves and assist with related statistical analysis. When plotting these curves using all 100k infringing extensions, the confidence intervals became too narrow and were not visible in the figure. Since this makes it difficult to assess the uncertainty around the survival estimates, we perform several experiments with different random sample sizes to find the best visualization. Seeing that the curves have no meaningful visible changes other than the confidence intervals, we take a random sample of 1,000 extensions (per variable or class) to display the Kaplan–Meier curves in a clearer manner. The curve labeled as “All” from Figure 5 shows the survival probability of infringing extensions. The median survival time is 581 days, confirming our earlier assertion that more than half of vetted extensions remain in the CWS for more than a year. However, while the distribution of vetted extensions shows few items with a life expectancy greater than a decade, the Kaplan–Meier estimate predicts that 6% of infringing extensions will surpass the same period. This result makes sense due to the inclusion of published infringing extensions as censored observations in the calculation of this curve.
We calculate multiple survival curves stratified by vetting label, and look for differences in the lifetime of vetted extensions based on this variable. Then, we perform log-rank tests between the three independent classes and obtain p-values smaller than 0.005 in all cases, thus rejecting the null hypothesis of no difference between groups. As such, we conclude that there is likely a significant difference in the lifetime of vetted extensions based on their vetting label. Figure 5 also shows separate Kaplan-Meier curves for these classes. We see substantial changes in the median survival time, with extensions flagged as policy violations having the longest expected lifetime (743 days), and malware having the shortest (282 days). The longer lifetime of policy violations may be related to the need for user complaints to initiate the process, although we do not have enough visibility into the vetting process to confirm this. Interestingly, policy violations tend to stay on the CWS for longer than minor policy violations (443 days), which is counterintuitive given the label name assigned by Google.
Takeaway. The CWS is excessively slow at removing infringing extensions, typically taking months or even years. At best, only 13% of extensions are vetted in less than a month. Malware is taken down faster than other types of infringing extensions, with a median lifetime of 282 days or well over half a year.
VI Characterizing Infringing Extensions
TABLE I: Top 65 largest infringing clusters. Includes only the first 15 NTE clusters, skipping the rest to provide a more representative sample. Footer row has totals from all clusters, including skipped. Highlighted rows contain published extensions. Extension labels are Total (T), Published (P), Policy Violation (PV), Minor Policy Violation (MPV), and Malware (M).
Cluster Example Extensions Publs. Users Lifetime
# Kind Extension ID Name T P PV MPV M Count Sum Avg Max
1 NTEs mngkjbegjgngjmbpojnmkngelecdodfm Fruits Basket New Tab… 3.6k 0 1.5k 2.1k 1 520 153k 1y 4y
2 NTEs cpcnejdpjegmjjljgfjklloeihdkdgbp Island Wallpapers The… 2.5k 0 829 899 822 258 181k 1y 1y
3 NTEs cbhjljcgicdehmecljnoggonepjpflij Tuzki New Tab Page HD… 2.5k 0 17 2.4k 0 253 42k 1y 1y
4 NTEs nidmpdeffdnhnccmmkfaagfoefhajmfl Lexus Wallpapers HD T… 1.8k 0 698 735 255 134 82k 1y 1y
5 NTEs meppkkplebidmpambckjodfhgheipfdp Superhero.io HD Wallp… 1.7k 3 69 1.6k 0 181 121k 1y 3y
6 NTEs fhidgmalgieecohkfdmehekjejlkiblo 4K Wallpaper HD Custo… 1.6k 0 33 573 9 145 18M 6m 1y
7 NTEs gdjpneilpaakfdkbofbdcoggimplfjek Dragon Ball Super Bro… 1.5k 0 240 1.3k 0 132 15k 2y 3y
8 NTEs gfdbfnafbpicmggajcjkehmdcnoadigh Sports car HD wallpap… 1.5k 0 149 1.4k 0 156 33k 5m 6m
9 NTEs lohlomonaokoijibhiaofphjkhpdmhhd Evan Fournier Themes … 1.5k 0 168 1.0k 0 144 78k 1y 4y
10 Unreliable dmhapbkbkeopdeapenongacbpajdfljg Calc LS 1.3k 53 49 1.1k 2 292 1M 4m 10y
11 NTEs jkocidindgkhipkkiaffdmoonpdpjkoo Bridgerton Wallpapers… 1.3k 0 2 783 0 80 174k 2m 4m
12 NTEs enihnbifjcjijpacpllggccehdlcomco My New Tab extension … 1.2k 0 0 1.2k 2 91 154M 1m 2m
13 NTEs emkadehobopegnnodjgmifldohhhehae Cityscape - City Wall… 1.2k 1 7 1.1k 0 79 5M 8m 1y
14 NTEs ngeclbmgabdohlkcpdjbgpoplghdhgbn Will Smith Popular St… 1.2k 0 669 491 0 199 46k 2y 3y
15 NTEs ckkplkjpdkpbnlcbbhldphbfmgcdfgcn Smite Wallpapers New … 1.1k 1 199 923 0 100 123k 2y 3y
16 Spam fpfieneffjbipoalhjkbjdeophjfgmhj WhatsApp Group Links 1.0k 67 59 740 1 300 545k 8m 3y
17 NTEs jafpgkdldemnlbmmcehiecehpigjploo Rwanda New Tab Rwanda… 905 0 297 585 2 76 54k 1y 1y
19 Kiosk nllpcefijadedeobnimfhahfpacimcgk Endless Runner 878 717 25 77 0 612 11k 5y 9y
24 Games afbcbdeiiofjclcghcjiackbcnebhlhi Umbrella Down 508 0 3 505 0 44 10k 1y 2y
25 Games gaiajadlfkefhennokenpgbflgmfelej WorldCraft: 3D Build … 501 0 4 493 0 22 29k 1m 2y
29 Apps cgpdaffgnmiapcmnmjlccgchnlnddhch Suecalandia 423 370 1 31 0 377 2M 8y 11y
33 Test nlmjjcpjklaapbblhokbnfbjahjodfkh Erase From View For T… 372 67 7 35 6 270 1.2k 1m 7y
34 Games mbmibbpjgfmhodpjhgcollomgggpkpbn Balls Avoid Arcade Ga… 316 300 10 6 0 45 1M 1y 1y
35 Replacer lijiloghjiehhmggfgloglklmbnoenen Impeachment Pie 314 198 11 95 0 300 11k 5y 8y
47 Unreliable niiicdjkhpheppjjaankjoegbbejdmho Magento Development C… 266 38 11 197 0 237 64k 3y 5y
48 Games himbonlbdfgjpnfdabgkomgghffbfhph Tarot Kariyer Falı Ba… 264 0 1 254 2 80 13k 5m 1y
49 APKs mdmkciaalenjmmannlibndcgjjiibdng ce 263 242 1 4 0 196 72k 7y 9y
54 Games gnpikgcjfmlcjpcnokpiindllmdmadem Digimon Battle Spirit… 229 0 1 227 1 22 23k 7m 3y
60 Boilerplate aemnapldoeccnbmfmifibgdjimkdgnig League of Legends Acc… 210 76 16 108 0 160 28k 4y 5y
63 Spam bbcfdgkbbllgjidioakddokoakeecloj Group Buy Seo Tools 206 189 0 11 0 125 1.6k 11m 2y
65 Games kjdcdphljphkdkipfhijimoedhpeheop Epic Charlie 201 0 2 199 0 14 4.3k 1y 3y
44k 2.4k 6.4k 29k 1.6k 6.2k 218M 1y 11y
This section examines what kinds of infringing extensions are the most prevalent in our dataset (RQ2). For this analysis, we take the infringing clusters obtained in Section V and sort them by descending size, prioritizing those with the most extensions. That is, we are interesting in finding the most common groups of behavior across infringing extensions. We select the top largest infringing clusters and assign a kind or type of content to each one by manually inspecting a random sample of the extensions it contains.
Table I shows the ranked list of the top 65 infringing clusters. Almost all of these clusters contain NTEs. To provide a broader picture of what kind of extensions lie in these clusters, we only display up to 15 NTE clusters in Table I, allowing for other kinds of clusters to appear. Therefore, rows that are missing from the table, such as cluster #18, are omitted because they also contain NTEs.
Top infringing clusters range in size from hundreds to a few thousand extensions. The number of unique publishers per cluster is always lower than its size, confirming that some developers republish the same extension with the same account. Out of the top 65 infringing clusters, containing 44k extensions, we only find 6.2k publishers. With some permissible exceptions allowed by the CWS, this is a strong indication of repetitive content [15]. NTE clusters have the highest removal rate, with almost no extensions still published at the end of our crawl. Interestingly, we find some instances of no consensus over the vetting label among extensions in the same cluster. For example, clusters #2 and #4 contain almost identical extensions that have been removed completely from the CWS, yet the reason varies among Minor Policy Violation, Policy Violation, and Malware. There is a high variance in the cluster impact, finding some with millions of affected users. The most severe case is cluster #12, which lasted just 2 months in the store but managed to gather a total of over 150 million users. In terms of lifetime, top infringing extensions last an average of one year. Some clusters like #34 are even fairly new, having been created a year ago and containing infringing extensions that have not yet been removed from the CWS.
We next characterize the most significant kinds of clusters we find in our results, providing examples for some of the extensions they contain.
VI-A New Tab Extensions (NTEs)
NTEs are, by far, the most repeated kind of extension, accounting for 83% of the top infringing clusters, or 37k extensions alone. NTEs override the default page that loads whenever a new tab is opened in the browser, offering more customization options and features. The common trait of NTEs is that they display a random background image every time this page loads. This has contributed to the proliferation of this type of extension in the CWS, as each NTE focuses on a particular theme or topic upon which the background images are based. Topics cover virtually every user interest imaginable, from abstract (colors, shapes), miscellaneous (cute animals, cars), places (Tokyo, New York skyline), sports, celebrities (people, music bands) or works of fiction (movies, video games).
To prevent publishers from flooding the CWS with low-quality, repetitive extensions, Google released new program policies in June 2021 that prohibited “multiple extensions with highly similar functionality,” specifically citing “wallpaper extensions” as an example [57]. Consequently, close to all NTEs found in the top infringing clusters have already been removed. We hypothesize that this high removal rate, in contrast to non-NTE clusters, is due to the fact that NTEs are easily findable because they include recurring keywords such as “wallpaper” or “new tab” in their names, facilitating the vetting process. Nevertheless, we still find 55 published NTEs in top infringing clusters, and at least 571 in the entire store (i.e., including clusters outside this ranking).
(a) Mountain Wallpaper [5]
(b) Salads Wallpaper [64]
Figure 6: Sample screenshots of published NTEs.
Apart from random background images, NTEs often provide other side features to stand out from other competing publishers. As shown in Figure 6, these features include a navigation bar with quick links to social media and other popular websites, shortcuts to the list of installed apps, bookmarks and most visited sites, draggable sticky notes, realtime weather forecast, and a search bar. Some extensions go a step beyond and override the default browser search engine, intercepting all searches typed in the omnibar.
While seemingly innocuous, NTEs may pose a risk to users due to the amount of personal information they have access to for legitimate purposes. Overriding the new tab page can be used to track browsing habits, whereas the list of top visited sites and searches help profile user interests. For this reason, we believe that clusters #2 and #4 contain hundreds of NTEs labeled as Malware instead of the customary Policy Violation. However, after manually examining a random subset, we find no meaningful behavioral differences between malware-labeled extensions and the rest of items in NTE clusters. As such, we conclude that CWS’s vetting labels are somewhat inconsistent and may differ between extensions with exactly the same set of behaviors or even codebase.
VI-B No Content Extensions
For a extension to be considered eligible for submission to the CWS, it must have a name, description, version, and an icon declared in its manifest [20]. Despite satisfying all the minimum requirements, an extension with only these attributes would be useless due to its lack of functionality. Furthermore, according to the Program Policies, it would violate the Minimum Functionality clause by providing no utility whatsoever to the CWS catalog [18].
Test Extensions. Cluster #33 contains extensions with the bare minimum manifest and an icon, lacking any HTML, CSS or JavaScript code. The items in this cluster date back as early as November 2015, with the most recent having been creating just a couple of days before the end of our crawl in March 2024. These extensions have names like “Free Trial Extension!” or “Search test,” and belong to publishers with dubious email addresses such as “cwsprodtest3@gmail.com” or that end at google.com, making the latter corporate Google accounts. We reported this cluster to the CWS Team asking for a clarification on the purpose of these extensions and whether they actually were uploaded by their employees. Unfortunately, we did not receive any response before the submission deadline.
Boilerplate Extensions. We find another cluster comprised of extensions with no functionality released between 2018 and 2022. These extensions have identical dummy background pages and a browser action that loads the example HTML document from Listing 1. Looking at this document, we notice that the extensions are made using Extensionizr, a now-defunct web application for generating boilerplate browser extensions in a few clicks [67]. While the extensions themselves do nothing, their store listings appear to promote blogs or businesses, with the featured image usually being a screenshot of the promoted website. As such, we posit these infringing extensions are just dummy items designed to take advantage of the CWS to advertise a given site.
<!doctype html>
<style type="text/css">
#mainPopup {
padding: 10px;
height: 200px;
width: 400px;
font-family: Helvetica, Ubuntu, Arial, sans-serif;
h1 {
font-size: 2em;
</style>
<div id="mainPopup">
<h1>Hello extensionizr!</h1>
[…]
</div>>>>>>>>
Listing 1: Default browser action document generated by Extensionizr.
VI-C Spam Extensions
Among the top infringing clusters, we find extensions whose sole purpose is to promote or spam a particular website or link. Much like no content extensions, these items provide the user with no apparent meaningful functionality. However, unlike the previous group, they do contain source code and assets, albeit with little effort. The implementation varies slightly by cluster, with all extensions in a cluster reusing the same manifest, file structure and much of their source code.
One example is cluster #16, which consists of extensions only declaring a browser action in their manifest. When clicking the extension icon, it loads an HTML document that has a link to the spammed website. These extensions rely on popular apps and services to grab the attention of CWS users, with names such as “123movie - 100% WORKING”, “Free Dogecoin Faucet” and “Whatsapp Plus APK”. Notably, “GS Auto Clicker:Free Download” is the only extension labeled as Malware in this cluster,444 https://crxcavator.io/report/chjhnkfpbgcajkfidohljkjjlfcmnahi which differs from the rest only in the link it promotes. Cluster #63 is also made up of extensions that have a browser action, but they take it a step further to give the appearance of having some functionality. This is achieved by having an HTML form with a few fields that relies on a simple JavaScript code to produce some—usually meaningless—output. This form also contains a call-to-action button promoting the spammed website. For example, the extension “arsenal vs everton live” has a football match score predictor that gives a random pair of numbers every time the form is submitted [35].
Games. A predominant type of content used to promote websites across top infringing clusters is games. These extensions generally follow the same implementation pattern observed in other spam clusters. Extensions in clusters #24 and #25 implement a browser action with a “Play” or “Play Now” button that opens the purported URL for the game in a new tab. Similarly, clusters #48 and #54 have a background page to open this URL in a new window when the extension icon is clicked. We also find publishers that make an effort to comply with the Minimum Functionality clause by embedding the actual game inside the extension instead of redirecting the user to an external website. Clusters #34 and #65 are instances of this case, containing Unity WebGL and Adobe Flash games, respectively. These extensions also have links to websites hosting games, with the distinction that they can work offline.
Unreliable Clusters. We find 2 other clusters resembling those containing spam extensions, as they also have very similar manifests that just declare a browser action. While these clusters are mostly comprised of extensions with no functionality and links to external websites, they also contain low-effort extensions that perform an extremely simple task (e.g., a calculator). Unlike spam extensions, items in unreliable clusters lack a distinct set of API calls that can be used to group them together, and so they end up in different clusters than the former. We acknowledge that this is a limitation of our pipeline as currently designed, since extensions with short manifests and no API calls are clustered together. We discuss this point further in Section VIII.
VI-D Other Extensions
In addition to the previous groups, we encounter other clusters of extensions that, while similar, are more difficult to classify as infringing content without further manual inspection.
Chrome Apps. Apps are a legacy type of extension that were deprecated in 2020 and are currently only supported on ChromeOS devices [11]. They can be as simple as a shortcut that opens a URL in a new tab, being essentially an icon and a manifest that declares the app.launch.web_url property. As such, neither our tool (nor any tool) can determine whether they violate any policies without analysis of the linked external site.
Kiosk Apps. Cluster #19 is formed by kiosk apps generated using the official Chrome App Builder extension [24]. Akin to Chrome Apps, these extensions just load an external URL, but are intended to run full-screen on a dedicated kiosk or signage device.
Repackaged APKs. ChromeOS devices can run Android apps through the App Runtime for Chrome (ARC) [12]. With this runtime, developers can bundle Android Application Package (APKs) into browser extensions and run them on a Chromebook. Because our pipeline cannot extract features from native APKs, we cannot cluster these extensions any further.
Text Replacers. We find a cluster of satirical extensions, often related to politics, with the sole purpose of replacing a pattern with another word or phrase on all visited webpages (e.g., “stocks” to “stonks” [sic]). They do so by declaring a content script in their manifest that injects a short JavaScript file to manipulate the DOM. We are unsure whether these extensions comply with the Minimum Functionality and Spam and Abuse policies. However, we note that a third of this cluster has been taken down from the CWS.
VI-E Labeling Consistency
As described in Section III, the CWS assigns a label to vetted extensions based on the reason for the takedown. After characterizing the top infringing clusters, we notice some discrepancies in the labels assigned by Google to these vetted extensions. This raises the question of whether the CWS vetting labels are indeed related to the behavior of an extension and if the process is consistent. To assess this, we identify which are the most common features among vetted extensions per label and compare the groups. We take all vetted extensions in our dataset, group them by vetting label, and count the occurrences of features in each group; that is, we count how many vetted extensions with a particular label have a given feature. We initially conducted this experiment including NTEs. However, doing so added too much noise to our results due to the large number of vetted NTEs and the seeming inconsistency of their vetting labels, as previously mentioned in Section VI-A. As such, we repeated the experiment and put extensions with “theme”, “wallpaper” or “new tab” in their name in a separate group.
TABLE II: Top most occurring features in vetted extensions grouped by label. NTEs reported separately, highlighted rows show features in any top 5.
Feature M PV MPV NTE
MV2 #1 60% #1 86% #1 93% #3 88%
storage #2 52% #2 45% #4 30% #8 46%
browser.tabs.create #3 51% #5 34% #3 30% #2 93%
browser.runtime.onInstalled #4 49% #8 22% #8 19% #4 87%
MV3 #5 40% #21 14% #31 7% #57 12%
browser.runtime.onMessage #6 40% #4 35% #5 21% #10 43%
browser.runtime.id #7 37% #30 11% #29 7% #7 47%
browser.storage.local.get #8 37% #12 20% #21 10% #23 22%
tabs #9 36% #3 41% #2 42% #20 23%
browser.storage.local.set #10 35% #13 19% #22 10% #22 22%
browser.tabs.query #13 31% #6 29% #9 18% #11 39%
browser.runtime.sendMessage #12 32% #7 27% #10 14% #13 36%
activeTab #53 7% #9 22% #7 21% #38 17%
navigator.userAgent #11 32% #10 21% #13 11% #59 10%
browser.browserAction.onClicked #22 18% #15 16% #6 21% #5 76%
chrome_url_overrides.newtab #34 11% #43 7% #28 7% #1 99%
browser.runtime.setUninstallURL #14 26% #36 9% #35 6% #6 57%
topSites #262 1% #129 2% #66 3% #9 45%
Table II shows the top most occurring features found in vetted extensions per vetting label. Overall, there is a considerable overlap between the feature sets of the reported groups, with 5 features consistently appearing in the top 10 rankings regardless of the vetting label. Furthermore, these repeated features are related to the storage, runtime and tabs extension APIs, which are widely used in published (non-vetted) extensions as well. Still, we notice some interesting aspects that are worth discussing.
According to Google, one of the key reasons for transitioning to Manifest V3 is its “higher security and privacy guarantees” [33]. We see that malware authors are already transitioning to it, with this group having the highest ratio of extensions using MV3 among the studied groups (40%). One of the top API calls used by extensions in the Minor Policy Violation group is browser.browserAction.onClicked. This method is used to add a listener that will run when the extension icon is clicked, typically to open a new tab or a browser action popup. This validates the evidence we discuss in Section VI-C. Regarding NTEs, we see that more than half of them declare a URL that will open in a new tab when the extension is uninstalled, a much higher ratio than for the other groups. We also find that the topSites permission is used almost exclusively by NTEs, making it a good indicator of this group across vetted extensions.
Based on these findings, we conclude that vetting labels are not strictly related to the behavior of an extension, especially in the case of NTEs, which appear in all three labels.
Takeaway. NTEs are the most common kind of infringing extension, accounting for 83% of the top 65 largest clusters. Extensions with no functionality that clearly do not comply with the store policies number in the thousands, with hundreds of items still being published. CWS vetting labels are somewhat inconsistent, not being strictly related to the behavior of an extension.
VII Characterizing CWS Malware
Out of all vetted extensions, those labeled as malware are of significant importance. In this section, we study which malware families are published in the CWS and use SimExt to cluster them by their behavior (RQ3). We conclude with an analysis of malicious extensions with a high number of installs that are not present in VirusTotal. We note that we do not claim to study the entire browser extension malware ecosystem, as some families may be distributed outside of the CWS by other means, such as sideloading [62, 68, 55].
VII-A Malware Detection
As mentioned throughout this paper, the CWS assigns its own labels, including a generic one for malware, to vetted extensions as part of the vetting process. As such, we ask ourselves how aware the threat intelligence community is about these extensions that the CWS considers malicious. To answer this question, we take all vetted extensions in our dataset that are labeled as Malware, compute the SHA-256 hashes of their CRX packages, and query VirusTotal to obtain labels from third-party security vendors. Based on these results, we group our subset of malware extensions into three categories: $(i)$ Not Found if the sample has not been uploaded to VirusTotal, $(ii)$ Clean for samples with zero detections, and $(iii)$ Malicious if it is flagged as malicious or suspicious by at least one engine.
Of the 5,647 extensions labeled by the CWS as Malware, a remarkable 95% of them (5,377 extensions) were not previously seen by VirusTotal; 3% of malware extensions fall into the Clean category; and merely 83 extensions (1%) are detected as Malicious. Having such a significantly high number of Not Found and Clean items for extensions that the CWS itself reports as malware raises some concerns. One implication is that most malware families that target the CWS are unknown to the threat intelligence community, possibly because these samples are just removed from the CWS (and from infected victims) but they are not shared with the community.
VII-B Malware Families
We group the few known malicious samples that we found by their most popular threat classification, or family, as suggested by VirusTotal. For each group, we count the number of extensions, measure the impact by the number of user installs, the coverage of the detection engines by the ratio of detections, and the lifetime of a family by the release dates of the first and last seen extension versions.
TABLE III: Families of malware extensions detected by engines in VirusTotal.
Exts. Detections Users Lifetime
Count Max Avg Sum From To
N/A 39 39% 10% 9M 2018-01-31 +5y
trojan. 28 17% 13% 59M 2020-05-08 +2y
browext 6 17% 13% 9M 2022-02-23 +1y
trojan.browext/chromex 4 21% 21% 365k 2022-08-26 +1m
trojan.chromex 3 32% 25% 2M 2022-01-19 +10m
adware.broextension 1 3% 3% 700k 2023-10-23 +1d
pua.keylogger/chromelogger 1 33% 33% 50k 2020-03-01 +1d
trojan.freesub/chromex 1 37% 37% 8.0k 2020-11-23 +1d
Table III shows the aforementioned groups of extensions that are known to VirusTotal per family. Almost half of the malicious extensions have no suggested threat classification label, closely followed by a generic “Trojan” group. Other labels such as “BrowExt”, “BroExtension” or “ChromeX” seem to only indicate that the malware is a browser extension, without adding any further information about the malware family or behavior. The only three extensions with more informative classifications are “SaveProtect VPN”,555 https://crxcavator.io/report/dodnpoijjkmcmlhlelmggejhfocfjgfc which is labeled as Adware; “Fea KeyLogger”,666 https://crxcavator.io/report/fgkghpghjcbfcflhoklkcincndlpobja which advertises itself as a keylogger; and “Free YouTube Subscribers Generator”,777 https://crxcavator.io/report/fdfchfidjajpidpjilnlboncflgnjdda which has a link to an external website instead of having any JavaScript code. From these findings, we conclude that vendor labels for malicious browser extensions are extremely poor, and in most cases non-existent.
Grouping only these known malicious extensions by behavior using SimExt, we end up with 7 different similarity clusters and 68 extensions that are classified as outliers. The resulting clusters have extensions that fall into the following types: $(i)$ media file downloaders, $(ii)$ cursor icon customizers, $(iii)$ volume boosters, $(iv)$ reader mode extensions, and $(v)$ ad blockers. Appendix B lists these extensions for more information.
VII-C Case Studies
Since almost all malware extensions flagged by the CWS are unknown to detection engines, we manually inspect those not present in VirusTotal and with at least 100k installs at the time of removal. Of the resulting 90 items that meet these criteria, we could not find a clear motivation for labeling 29 of them as malware.
A large group of the remaining extensions contain extensions with embedded tracking capabilities or that load remotely hosted code. One interesting group is formed by 12 extensions that use Google Tag Manager to download and execute an obfuscated script888 https://storage.googleapis.com/g1analytics/cloud_new_noab-obf.js from a Google Cloud Storage bucket. When unpacked, this script sends an HTTP request to a remote endpoint to get the country of the user and starts exfiltrating all visited URLs in real time if the location matches the United States. To avoid detection, the payload of these requests is encoded as a binary blob. Some extensions that request the webRequestBlocking permission load an additional remote script,999 https://storage.googleapis.com/analytics-cloud/js_analytics_protected.js which intercepts all search queries from a predefined list of lesser known search engines and redirects them to another domain.
We also find 4 trojanized extensions that run code in Facebook to obtain authentication tokens (such as fb_dtsg). The intercepted tokens are then used to make requests to the GraphQL endpoint in the background without the user’s awareness. Once authenticated, the user is joined to a Facebook group that varies depending on the extension, and the device’s access token is posted there. Along with the token, these extensions also exfiltrate the number of business accounts associated with that credential, presumably to triage victims.
Appendix C lists the extensions discussed above.
Takeaway. Only 1% of extensions labeled as malware by the CWS are considered malicious by third-party security vendors integrated on VirusTotal. Furthermore, these vendors rarely assign a malware family to these samples, and when they do, they use generic labels that provide little information.
VIII Discussion
This section discusses the key findings of our work and provide recommendations that could help improve the CWS vetting process. We also discuss the limitations of our analysis, particularly of SimExt, and describe future work to address them.
Detection of Infringing Extensions. The findings outlined in Section V support the assumption that the current CWS vetting process lacks scalable tooling for effectively finding or clustering similar extensions. Given the estimate that 86% of infringing extensions found by our tool are republished extensions (i.e., items that the store has seen and taken down before), we recommend using not only blocklists but also behavioral features to flag similar extensions, as this can contribute to improve the vetting process. We believe that similarity search tools such as SimExt can be a valuable complement to the human factor, helping to find infringing content faster and more accurately.
Repeat Offenders. We find evidence that the CWS does not ban publishers with multiple vetted extensions. Since 11% of the published infringing extensions found by our tool come from repeat offenders, we believe the CWS should enforce their own Repeat Abuse policy [19] and suspend the accounts of repeat offenders.
Lifetime of Infringing Extensions. Infringing content stays for too long in the store, with 59% of vetted extensions remaining published for more than a year before they are taken down. Even for malware extensions, the median removal time is longer than 9 months at best. We consider these to be unacceptably high and believe that the CWS needs to focus its efforts on significantly reducing them. The addition of automated tools for finding items similar to previously vetted extensions should also help reduce these lifetimes.
Unpublished Infringing Extensions. As presented in Section V, we find that some publishers of infringing extensions voluntarily remove them before receiving a takedown from Google. Given that repeat offenders are fairly common (and thus there is little repercussion for having multiple vetted extensions), we do not believe that these developers are acting like this to protect their accounts. Extensions labeled by Google as malware are remotely and automatically uninstalled from users browsers [7]. However, we find that extensions that are unpublished before being detected as malware are not subject to this process, leaving users compromised. Thus, we recommend that the CWS also assigns vetting labels to recently unpublished extensions as well as to vetted ones.
Vetting Labels. Providing labels for vetted content is a valuable transparency feature of the CWS vetting process. However, the current labeling scheme admits multiple improvements. One would be to move towards finer-grained labels that are more specific about the reasons for taking down the item, as these labels could better inform ecosystem studies such as ours. As mentioned in Section VI-A, more consistency when assigning labels to vetted extensions will also help increase confidence in the vetting process.
Malware Labels. Given that only 1% of extensions labeled as malware by the CWS are detected by security vendors in VirusTotal, we strongly recommend that the threat intelligence community improves malware detectors for browser extensions. More fine-grained family labels will greatly assist in classifying and tracking trends in the browser extension malware ecosystem.
VIII-A Limitations and Future Work
In Section VI-C, we discuss how our pipeline clusters together extensions with very few features, to the point that items with no logged API calls and different behaviors are difficult to distinguish from one another. Ideally, these unreliable clusters should be split further into as many sub-clusters as unique sets of behaviors. As an improvement, we propose expanding the API calls extracted by our static and dynamic analyses to include other frequently used objects, such as document and jQuery. This way, extensions with few or no calls to Extension APIs or inside the navigator object will be better clustered due to the use of more robust embeddings.
IX Related Work
Previous research has crawled the CWS for various purposes related to the analysis of browser extensions [29, 66, 1, 3, 48, 9, 31, 58, 8, 51, 44].
Extension Analysis. Pantelaios et al. statically extracted API calls to cluster extension version deltas to detect malicious updates [48]. Fass et al. used static analysis to find suspicious data flows in vulnerable extensions [9]. Kapravelos et al. proposed a dynamic analysis architecture for identifying malicious behavior in extensions using on-the-fly generated pages [29]. Jagpal et al. used both static and dynamic analysis to capture behavioral signals from browser extensions [28], followed by Aggarwal et al. a few years later [1]. Picazo-Sanchez et al. used static, manual and dynamic analysis to mark malicious extensions [51]. Eriksson et al. combined both static and dynamic analysis to discover code vulnerabilities, uncovering the existence of NTEs that stole browsing traffic [8]. Pantelaios and Kapravelos used dynamic execution to validate automated conversions of extensions from MV2 to MV3 [47]. Unlike our work, previous dynamic analysis efforts have relied on running the extensions on an actual web browser through manual interaction or using tools like Selenium, Puppeteer and Playwright. Instead, we created a mocked V8 sandbox for faster code execution at scale.
Defining Potentially Harmful Extensions. Somé was the first to propose a taxonomy of sensitive APIs to define extensions that pose a security or privacy risk to the user [59]. Hsu et al. introduced the concept of Security-Noteworthy Extensions to extend this set to cover malware, vulnerable, and policy-violating extensions [26]. In our work, we focus on evaluating the CWS vetting process by using its own vetting labels, rather than seeking to define what potentially harmful extensions are.
Finding Malicious Extensions with Machine Learning. Jagpal et al. used machine learning to flag malicious extensions based on behavioral signals, having to train a proprietary model daily to account for newly vetted extensions [28]. Aggarwal et al. improved upon this by feeding sequences of API calls to their own Recurrent Neural Network (RNN) [1]. Pantelaios et al. used DBSCAN to cluster custom-made API sequences referred to as seeds [48]. Similarly, Picazo-Sanchez et al. used time series analysis and machine learning to cluster malicious extensions based on their download patterns [51]. In contrast to previous work, our approach does not try to classify extensions as benign or malicious, but to cluster similar extensions together. By using NLP and ZSL, we avoid the limitation of having to retrain our model when a new malicious behavior is discovered.
To the best of our knowledge, no prior work has developed a comprehensive methodology for measuring similarity between extension based on static features and dynamic behavior, nor used vector embeddings to cluster browser extensions.
X Conclusion
This paper has presented a comprehensive study of the CWS vetting process and the prevalence of infringing content. To assist in this analysis, we developed SimExt, a novel methodology for measuring similarity among browser extensions. Our tool has proven to be instrumental to find infringing content that otherwise went unnoticed. We also believe that SimExt may be valuable in other application domains. Informed by our findings, we present several recommendations that could contribute to improving the vetting process and the analysis of malicious and infringing content.
Appendix A Serialization Example
Figure 7 provides an example of the feature serialization pipeline we describe in Section IV-B. The process takes as input the manifest key-value pairs of the extension and the API calls extracted by the static and dynamic analyzers. These items are converted to sentences that can then be merged into a single document.
Figure 7: Feature serialization example.
Appendix B Malicious Clustered Extensions
Table IV lists extensions detected as Malicious by third-party security vendors that belong to a similarity cluster obtained using SimExt.
TABLE IV: Known malicious extensions that belong to a similarity cluster.
Extension Name
Media downloaders
pocmgnhmjgjghodelfkhbjaoidmbadpo Spotify™ & Deezer™ Music …
nhemnekeahfdfemcchogmiinhkpdbedp Spotify™ & Deezer™ Music …
hpbohmeoofibpbiiklpofdfehodejbmk VLC Video Downloader
dhnkeanajeheaifbmgalejfoebggoggk VLC Video Downloader
nadenkhojomjfdcppbhhncbfakfjiabp Base Image Downloader
okclicinnbnfkgchommiamjnkjcibfid Tap Image Downloader
Cursor customizers
magnkhldhhgdlhikeighmhlhonpmlolk Craft Cursors
pbdpfhmbdldfoioggnphkiocpidecmbp Clickish fun cursors
hdgdghnfcappcodemanhafioghjhlbpb Cursor-A custom cursor
Volume boosters
chmfnmjfghjpdamlofhlonnnnokkpbao Soundboost
hinhmojdkodmficpockledafoeodokmc HyperVolume
Reader mode extensions
icnekagcncdgpdnpoecofjinkplbnocm Easyview Reader view
dppnhoaonckcimpejpjodcdoenfjleme Readl Reader mode
Ad blockers
obeokabcpoilgegepbhlcleanmpgkhcp Venus Adblock
bkpdalonclochcahhipekbnedhklcdnp Epsilon Ad blocker
Appendix C Malware Extensions From Case Studies
Table V contains relevant items from the top most popular extensions labeled as malware by the CWS that do not appear in VirusTotal.
TABLE V: Relevant popular malware extensions that do not appear in VirusTotal.
Extension Name
Facebook stealers
ffmdedmghpoipeldijkdlcckdpempkdi Bookmarks Menu
dkpedpjjafnceedhomeijlphmjbblmdj Currency Converter PRO
jmphljmgnagblkombahigniilhnbadca Open link in same tab, pop-up …
adkpffmlkncmmimpnmogphiijidakdhm bilibili
Google Tag Manager
lfagjcmdalpklemkmdcblfghhkjjohbm Colorize Facebook
gahgachhcblgfnjdfghcjcpgbkbadfgg Easy Font Changer
lkcdbmaggddpfmfdbcloicogiaoepddk Floating video plus
njkmonlnhpfkaldcnhikggmdaepedcep Instant Eyedropper
jcjhgomglcabcikgghokgnheeeobakkb La notte
cnfianechkepmfdoakelcbamnbfbecke Loudly
idgifckkbacpebckkblhaopkfeikgipf Oscura dark theme
eemiojeoeomfggoapmnfnmpnkieojonj PDF tools all-in-one
konkphcpahjcebjdfkeihbalppeicalp Top Video Downloader
ailljajgcdcaadgmbncpfnofjanoabfn Video Download Center
kdnlfofefaichijbmflgibbdlfdapmbe Youtube Color Changer
mgccclinjajhpeiciiaflagddlhcillp Zoom it